# Yu Chen, Shuai Zheng, Nianyi Wang, Menglong Jin, Yan Chang Corresponding author

### v1

@@ Abstract
Simulating fuel sloshing within aircraft tanks during flight is crucial for aircraft safety research. Traditional methods based on Navier-Stokes equations are computationally expensive. In this paper, we treat fluid motion as point cloud transformation and propose the first neural network method specifically designed for simulating fuel sloshing in aircraft. This model is also the deep learning model that is the first to be capable of stably modeling fluid particle dynamics in such complex scenarios. Our triangle feature fusion design achieves an optimal balance among fluid dynamics modeling, momentum conservation constraints, and global stability control. Additionally, we constructed the Fueltank dataset, the first dataset for aircraft fuel surface sloshing. It comprises 320,000 frames across four typical tank types and covers a wide range of flight maneuvers, including multi-directional rotations. We conducted comprehensive experiments on both our dataset and the take-off scenario of the aircraft. Compared to existing neural network-based fluid simulation algorithms, we significantly enhanced accuracy while maintaining high computational speed. Compared to traditional SPH methods, our speed improved approximately 10 times. Furthermore, compared to traditional fluid simulation software such as Flow3D, our computation speed increased by more than 300 times.
@@ Introduction
Aircraft fuel sloshing simulation intersects multiple disciplines, including computer graphics, 3D computer vision, fluid dynamics, etc. The fuel tank is a key component of an aircraft fuel system. During large-angle maneuvers, fuel sloshing can cause violent fluid deformation and shifts in the center of gravity, both of which can endanger flight safety. Consequently, the simulation of fuel sloshing during flight is a critical research topic in aviation applications.
There are two typical representations of fluid in the area of Computational Fluid Dynamics: grid-based methods and particle-based (also known as SPH). Traditional fuel sloshing simulations typically use grid-based fluid simulation software such as Ansys Fluent and Flow3D. The grid-based approach segments the fluid region into a grid of cells or voxels. This approach can maintain high numerical stability and has been widely applied in engineering. However, it requires significant computation time and storage, often taking weeks or even months to complete, and it struggles with handling complex boundaries. In recent years, fluid simulation methods based on SPH (Solenthaler and Pajarola 2009; Bender and Koschier 2015; Ye et al. 2019; Li, Liu, and Zheng 2018; Koschier et al. 2020) have gained traction in fuel sloshing simulation (Zheng et al. 2021; Calderon Sanchez et al. 2019), such as DualSPHysics (Alshaer, Rogers, and Li 2017; Pourabdian, Omidvar, and Morad 2017), due to their relatively higher efficiency and better handling of free surface flows. However, despite these advantages over grid-based methods, the inherently high computational cost of solving Navier-Stokes equations remains, which is highly inconvenient for practical applications.
@@
However, existing neural network methods (Ummenhofer et al. 2019; Shao, Loy, and Dai 2022; Prantl et al. 2022; Li et al. 2018) are not sufficiently robust, which is suitable only for simple scenarios, such as the freefall of fluid blocks or dam-break flows. In contrast, the internal structure of a fuel tank is much more complex, with ribs and holes that allow fuel flow between compartments. Current neural network methods lack sufficient learning capacity and the necessary constraints to guide the modeling of fluid dynamics in such intricate scenarios, making them prone to breakdown under such challenging conditions.
@@
    We propose a neural network that is the first to be robust enough to simulate fuel sloshing in complex fuel tanks, achieving a speed improvement of over 300 times compared to traditional fluid simulation software such as Flow3D, and nearly 10 times compared to traditional SPH methods.
@@
    We have constructed the first fuel sloshing surface dataset, encompassing 320,000 frames across four typical tank types and covering various possible flight maneuvers. Compared to existing datasets, this dataset features more complex fluid motions and scenarios, including multi-directional rotations.
@@
    We conducted comprehensive experiments on the Fueltank dataset and aircraft takeoff generalization scenarios, demonstrating that our network surpasses all previous methods in multiple evaluation metrics.
@@
Besides the local interactions between fluid particles, we also note that learning global fluid features can enhance the neural network’s performance in global stability control. Previous works can effectively extract global features from point clouds (Qi et al. 2017a, b; Chen and Shi 2023; Wang et al. 2019; Wu et al. 2024). In this task, we require a simple and efficient global feature extraction layer, which enables the neural network to broadly control the overall fluid motion within a reasonable range (e.g., fluid particles in a closed box should not move outside the box). The extracted global feature does not need to be highly detailed, as fluid dynamics primarily focuses on the interactions between local particles. The global feature serves as additional information to make the framework more robust.

### v3

# A Pioneering Neural Network Method for Efficient and Robust Fluid Simulation

@@ Abstract
Fluid simulation is an important research topic in computer graphics (CG) and animation in video games. Traditional methods based on Navier-Stokes equations are computationally expensive. In this paper, we treat fluid motion as point cloud transformation and propose the first neural network method specifically designed for efficient and robust fluid simulation in complex environments. This model is also the deep learning model that is the first to be capable of stably modeling fluid particle dynamics in such complex scenarios. Our triangle feature fusion design achieves an optimal balance among fluid dynamics modeling, momentum conservation constraints, and global stability control. We conducted comprehensive experiments on datasets. Compared to existing neural network-based fluid simulation algorithms, we significantly enhanced accuracy while maintaining high computational speed. Compared to traditional SPH methods, our speed improved approximately 10 times. Furthermore, compared to traditional fluid simulation software such as Flow3D, our computation speed increased by more than 300 times.
@@ Introduction
Fluid simulation intersects multiple disciplines, including computer graphics, 3D computer vision, fluid dynamics, etc. There are two typical representations of fluid in the area of Computational Fluid Dynamics: grid-based methods and particle-based (also known as SPH). Traditional fluid simulations typically use grid-based fluid simulation software such as Ansys Fluent and Flow3D. The grid-based approach segments the fluid region into a grid of cells or voxels. This approach can maintain high numerical stability and has been widely applied in engineering. However, it requires significant computation time and storage, often taking weeks or even months to complete, and it struggles with handling complex boundaries. In recent years, fluid simulation methods based on SPH (Solenthaler and Pajarola 2009; Bender and Koschier 2015; Ye et al. 2019; Li, Liu, and Zheng 2018; Koschier et al. 2020) have gained traction in fluid simulation (Zheng et al. 2021; Calderon Sanchez et al. 2019), such as DualSPHysics (Alshaer, Rogers, and Li 2017; Pourabdian, Omidvar, and Morad 2017), due to their relatively higher efficiency and better handling of free surface flows. However, despite these advantages over grid-based methods, the inherently high computational cost of solving Navier-Stokes equations remains, which is highly inconvenient for practical applications.
@@
However, existing neural network methods (Ummenhofer et al. 2019; Shao, Loy, and Dai 2022; Prantl et al. 2022; Li et al. 2018) are not sufficiently robust, which is suitable only for simple scenarios, such as the freefall of fluid blocks or dam-break flows. In contrast, the internal structure of a tank with ribs and holes is much more complex. Current neural network methods lack sufficient learning capacity and the necessary constraints to guide the modeling of fluid dynamics in such intricate scenarios, making them prone to breakdown under such challenging conditions.
@@
    We propose a neural network that is the first to be robust enough to simulate fluid dynamics in complex tanks, achieving a speed improvement of over 300 times compared to traditional fluid simulation software such as Flow3D, and nearly 10 times compared to traditional SPH methods.
@@
    We have constructed the first complex fluid sloshing surface dataset, encompassing 320,000 frames across four typical tank types and covering various possible flight maneuvers. Compared to existing datasets, this dataset features more complex fluid motions and scenarios, including multi-directional rotations.
@@
    We conducted comprehensive experiments on the fluid dataset and aircraft takeoff generalization scenarios, demonstrating that our network surpasses all previous methods in multiple evaluation metrics.
@@
Besides the local interactions between fluid particles, we also note that learning global fluid features can enhance the neural network’s performance in global stability control. Previous works can effectively extract global features from point clouds (Qi et al. 2017a, b; Chen and Shi 2023; Wang et al. 2019; Wu et al. 2024; Liu et al. 2022). In this task, we require a simple and efficient global feature extraction layer, which enables the neural network to broadly control the overall fluid motion within a reasonable range (e.g., fluid particles in a closed box should not move outside the box). The extracted global feature does not need to be highly detailed, as fluid dynamics primarily focuses on the interactions between local particles. The global feature serves as additional information to make the framework more robust.